### v1

# Fantastic Pretraining Optimizers and Where to Find Them Kaiyue Wen

Abstract
AdamW has long been the dominant optimizer in language model pretraining, despite numerous claims that alternative optimizers offer 1.4 to 2 speedup. We posit that two methodological shortcomings have obscured fair comparisons and hindered practical adoption: (i) unequal hyperparameter tuning and (ii) limited or misleading evaluation setups.
To address these two issues, we conduct a systematic study of ten deep learning optimizers across four model scales (0.1B-1.2B parameters) and data-to-model ratios (1–8 the Chinchilla optimum). We find that fair and informative comparisons require rigorous hyperparameter tuning and evaluations across a range of model scales and data-to-model ratios, performed at the end of training.
First, optimal hyperparameters for one optimizer may be suboptimal for another, making blind hyperparameter transfer unfair.
Second, the actual speedup of many proposed optimizers over well-tuned baselines is lower than claimed and decreases with model size to only 1.1 for 1.2B parameter models. Thirdly, comparing intermediate checkpoints before reaching the target training budgets can be misleading, as rankings between two optimizers can flip during training due to learning rate decay.
Through our thorough investigation, we find that all the fastest optimizers such as Muon and Soap, use matrices as preconditioners — multiplying gradients with matrices rather than entry-wise scalars. However, the speedup of matrix-based optimizers is inversely proportional to model scale, decreasing from 1.4 over AdamW for 0.1B parameter models to merely 1.1 for 1.2B parameter models.
1 Introduction
Figure 1: Top Left:
The commonly used AdamW baseline for optimizer design is under-tuned. Up to a 2 speedup is achievable by tuning a single hyperparameter (learning rate) in the GPT-3 recipe Brown et al. (2020) for a 100M model (adopted in Liu et al. (2024a); Wen et al. (2024); Yuan et al. (2025); Liang et al. (2025); Wang et al. (2025)), highlighting the importance of proper hyperparameter optimization.
Top Right: Fixing hyperparameters across optimizers does not guarantee fair comparison. Shared hyperparameters such as learning rate and weight decay are commonly set to a constant in previous studies. However, even conceptually similar optimizers may correspond to very different optimal hyperparameters.
Bottom Left: Speedup decays with model size. While some optimizers show high (1.3-1.4) speedup over AdamW on models under 1B parameters, the speedup decays with model size to only 1.1 for 1.2B parameters.
Bottom Right: Matrix-based optimizers consistently outperform scalar-based optimizers.
The loss curves for three scalar-based optimizers (AdamW, Nesterov AdamW, Mars) and three matrix-based optimizers (Kron, Soap, Muon) trained with different Chinchilla ratios of data are shown. Matrix-based optimizers achieve a consistent speedup over scalar-based optimizers. Furthermore, the three matrix-based optimizers converge to a similar loss in an overtrained setting.
Pretraining has been the most computationally expensive component in the training pipeline for large language models, accounting for over 95% of the cost in DeepSeek V3 DeepSeek-AI et al. (2025b), and the additional RL training cost in DeepSeek R1 DeepSeek-AI et al. (2025a) is also comparatively much smaller.
Until recently, AdamW has been the standard optimizer. Recent studies have introduced novel optimizers that claim to accelerate pretraining by 1.4 to 2 compared to AdamW Liu et al. (2024a); Vyas et al. (2025); Liu et al. (2025a); Yuan et al. (2025); Liang et al. (2025); Wang et al. (2025); Liu et al. (2025c); Pethick et al. (2025); Ma et al. (2025), yet these optimizers have not yet been widely adopted in real-world pretraining DeepSeek-AI et al. (2024); Yang et al. (2025); Grattafiori et al. (2024), with the notable exception of Kimi K2 Team et al. (2025), which uses the Muon-clip optimizer (a variant of Muon Jordan et al. (2024)).
We pinpoint two problems in optimizer evaluation in the evaluation process of these optimizers that both undermine confidence in new methods and limit their practical adoption. First, some baselines suffer from improperly tuned hyperparameters. Second, many experiments are confined to smaller-scale settings, leaving unanswered questions about how these optimizers perform in broader, more realistic scenarios.
To address these issues, we benchmark eleven optimizers, including AdamW, in a rigorously controlled setup, focusing on two key questions:
1.
How to ensure hyperparameter optimality? Previous works typically rely on manual hyperparameter selection and keep common hyperparameters such as learning rate and weight decay fixed across optimizers (e.g. Liu et al. (2025a)). This tuning process may result in a weak baseline, as the hyperparameters chosen may favor the proposed optimizers rather than for AdamW. We validate this concern by showing that tuning just one hyperparameter in the widely adopted GPT-3 recipe (introduced in Brown et al. (2020) and used in Liu et al. (2024a); Wen et al. (2024); Yuan et al. (2025); Liu et al. (2025d, c); Liang et al. (2025); Wang et al. (2025)) can yield a 2 speedup for pretraining (Figure˜1, top left). To address, this we perform coordinate descent in the hyperparameter space on N hyperparameters for all eleven optimizers, iterating until convergence across for each of six different settings for models with up to 0.5B parameters to ensure that the hyperparameters are near-optimal for each setting.
2.
How do speedups differ across different scaling regimes? Large language models are trained in different regimes: depending on the model size, the data-to-model ratio (the number of tokens over the number of parameters) can range from to more than times the Chinchilla optimal (about based on Hoffmann et al. (2022)). Typical optimizer experiments in Chinchilla optimal regimes raise concerns about a new optimizer’s effectiveness in high data-to-model ratio. To address this, we benchmark the optimizers in four distinct data-to-model ratios (1, 2, 4 and 8 the Chinchilla optimal regime) and scale up to 1.2B parameter models (following Everett et al. (2024); Zhang et al. (2025a)).
Concretely, we employ the Llama 2 architecture Touvron et al. (2023b); Grattafiori et al. (2024)
(ranging from 0.1B to 1.2B parameters) and a data mixture similar to OLMo 2’s OLMo et al. (2025). We focus on the final validation loss on the C4-EN mixture as a known proxy for downstream performance Bhagia et al. (2024), while also tracking exact downstream performance on various benchmarks. As previous works Vyas et al. (2025); Liu et al. (2025a) show that the step-wise computation overhead of matrix-based optimizers can be reduced to under through proper implementation, we primarily compare algorithms by the number of tokens needed to reach a given loss.
Our empirical results show the necessity of careful hyperparameter tuning and end-of-training evaluations across a range of model scales and data-to-model ratios:
wo
1.
Hyperparameter transfer between optimizers is non-trivial. Even similar optimizers may need very different hyperparameters (e.g., Lion’s optimal weight decay 0.6 vs. AdamW’s 0.1, Figure˜1, top right), so fixing hyperparameters across optimizers can lead to unfair comparisons.
2.
The speedup of new optimizers is lower than claimed and diminishes with model size. Many reported speedups of 2 simply reflect a weak baseline. Against our well-tuned AdamW baseline, the speedup of alternative optimizers does not exceed 1.4(Figure˜3). Furthermore, while new optimizers such as Muon and Soap show 1.3 speedups for small models (0.1B), the speedups diminish to around 1.1 for 1.2B parameter models at 8 Chinchilla Ratio (Figure˜1, bottom left), a regime that is not tested in previous works studying the scaling law of these optimizers 111The original Soap paper Vyas et al. (2025) investigate model sizes up to 0.6B parameters. and Kimi’s paper on Muon Liu et al. (2025a) only considers Chinchilla regime..
3.
Early-stage loss curves can mislead significantly. During learning rate decay, loss curves of different optimizers may cross multiple times (Figure˜5), so judging optimizers using intermediate checkpoints may result in a different ranking than comparing models at the target training budget.
Our benchmarking also reveals new insights about optimizer design:
1.
Matrix-based optimizers consistently outperform scalar-based optimizers for small models.Scalar-based optimizers (e.g., AdamW, Lion, Mars, etc.) update each parameter individually using scalar operations. After proper tuning, all scalar-based optimizers achieve similar optimization speeds to AdamW, with an average speedup ratio of less than 1.2. Matrix-based optimizers (e.g., Kron, Muon, Soap, etc.) leverage the inherent matrix structure of neural network parameters and precondition gradients using matrix multiplication. Despite their diverse update rules, matrix-based optimizers all deliver approximately a 1.3 speedup over AdamW (Figure˜1, bottom right) for models under 520M parameters.
2.
Optimal choice of optimizer shifts depends on data-to-model ratios. A winner in the Chinchilla regime may be suboptimal when data-to-model ratio increases. For example, while Muon is consistently the best optimizer in smaller Chinchilla ratio regimes, it is outperformed by Kron and Soap when the data-to-model ratio increases to 8x or larger (Figures˜3 and 4).
2 Related Works
Optimizers for Deep Learning. A long line of work has studied optimization
for deep learning, incorporating insights from classical optimization literature Robbins and Monro (1951); Nesterov (1983); Duchi et al. (2011) and domain knowledge about deep neural networks Sutskever et al. (2013). (i) Early insights motivated the rise of optimizers that use adaptive learning rates based on second-order momentum Tieleman (2012); Zeiler (2012); Kingma and Ba (2017). Adam Kingma and Ba (2017) later became the default baseline for optimizers with adaptive learning rates. Since then, improvements over Adam and SGD have been proposed, with notable examples including Nesterov Adam Dozat (2016) and AdamW with decoupled weight decay Loshchilov and Hutter (2019). Other improvements include addressing the convergence of Adam on convex loss Reddi et al. (2018); Zaheer et al. (2018); Taniguchi et al. (2024), considering interpolation between Adam and SGD to improve generalization Luo et al. (2019); Xie et al. (2022), performing further variance reductions on optimizer updates Liu et al. (2021); Zhang et al. (2019); Yuan et al. (2025); Xie et al. (2024); Pagliardini et al. (2024); Zhuang et al. (2020), incorporating momentum on weights Ivgi et al. (2023); Defazio et al. (2024b), allowing easier hyperparameter tuning Defazio and Mishchenko (2023); Mishchenko and Defazio (2024); Defazio et al. (2024a), reducing memory usage by incorporating the structure of
neural networks Shazeer and Stern (2018); Zhang et al. (2025b); Zhu et al. (2025); Luo et al. (2023); Modoranu et al. (2024); Zhao et al. (2024), and modifying the algorithm to allow larger batch sizes You et al. (2017, 2020).
(ii) Starting from Preconditioned SGD Li (2018a) and Shampoo Gupta et al. (2018), another line of optimizer design Morwani et al. (2024); Eschenhagen et al. (2023); Martens and Grosse (2020); Li (2018b); Eschenhagen et al. (2025) began to incorporate matrix preconditioners instead of simple scalar preconditioners. Techniques including learning rate grafting Agarwal et al. (2020), blocking, and distributed methodology Anil et al. (2021) have since been proposed. These matrix-based approaches later led to the theory of modular duality in deep learning optimization Bernstein and Newhouse (2024a, b); Large et al. (2024) and new optimizers including Muon Jordan et al. (2024) and Scion Pethick et al. (2025).
(iii) Motivated by Newton’s algorithm, there has been a line of work that tries to incorporate Hessian information Becker and Cun (1989); Yao et al. (2021); Schaul et al. (2013); Yao et al. (2021). (iv) Symbolic discovery of optimizers Chen et al. (2023) has discovered a memory-efficient SignGD Bernstein et al. (2018) variant called Lion, which claims to outperform Adam on a wide range of tasks.
Optimization for Pretraining. Since Brown et al. (2020), the cost of pretraining has increased dramatically. One of the challenges is how to choose hyperparameters with minimal cost.
A pivotal line of work in this direction is the tensor program series that allows for extrapolating some hyperparameters across scales Yang (2020b, c, a, 2021); Yang and Littwin (2023); Yang et al. (2022, 2024). Empirical results suggest that fitting a power law to scale hyperparameters is now common practice for large models Li et al. (2025a); Everett et al. (2024); Liu et al. (2025a); DeepSeek-AI et al. (2024); Zhang et al. (2025a). We have also incorporated this approach in our paper.
Another popular line of research is designing better optimizers specifically for pretraining. These optimizers are our main objects of study. An (incomplete) list of optimizers and their claimed speedups over AdamW includes Sophia Liu et al. (2024a) (2), Soap Vyas et al. (2025) (1.4), Muon Jordan et al. (2024); Liu et al. (2025a) (2), MARS Yuan et al. (2025) (2), Cautious AdamW Liang et al. (2025) (2), Block-wise Learning Rate Adam Wang et al. (2025) (2), FOCUS Liu et al. (2025c) (2), SWAN Ma et al. (2025) (2), DION Ahn et al. (2025) (3), and SPlus Frans et al. (2025) (2). We present a comparison of setups with these prior works in Appendix˜G.
Re-evaluation Methodology. Our work is a rigorous evaluation of optimizers for pretraining. Rigorous evaluation has been an important part of deep learning research to clarify the current status of research and move the community forward. Jiang et al. (2019) critically examines metrics for predicting LLMs’ generalization capability and has facilitated research on understanding loss landscape sharpness and generalization and new optimizers such as SAM Foret et al. (2021). Schmidt et al. (2021) re-evaluated optimizers at that time and showed that (i) which optimizer is optimal is problem-specific, and (ii) rigorous hyperparameter tuning is required and unequal tuning can account for most of the claimed speedup. Unlike Schmidt et al. (2021), we evaluate the LLM pretraining task and include modern optimizers they did not test. However, we arrive at a similar conclusion: rigorous and fair hyperparameter tuning is still not the norm, but rather the exception in optimizer design research. Zhao et al. (2025) also examines how different optimizers perform on the pretraining tasks. However, the focus of Zhao et al. (2025) is on understanding loss structure, and the optimizers tested in the paper are shown to have slower convergence compared to AdamW, whereas the optimizers tested in this paper show small but significant improvements in convergence speed. The Algoperf competition Kasimbeg et al. (2025) evaluates different optimizers across different settings and arrives at a similar conclusion that (i) matrix-based optimizers and (ii) denoising methods such as Nesterov momentum lead to speedup over AdamW. Our works focus on the pretraining setting and investigates the effect of scaling dataset and model sizes.
3 Methodology
In this section, we detail the experimental design and evaluation protocol that underpin our empirical investigation. In Section˜3.1, we specify the general setup for all subsequent studies. We then describe our three-phase hyperparameter-tuning framework: Phase I (in Section˜3.2) performs fine-grained coordinate-descent sweeps across multiple model sizes and data-to-model ratios to identify scaling-sensitive parameters; Phase II (in Section˜3.3) refines these sensitive parameters on mid-scale settings and selects the most promising optimizers; and Phase III (in Section˜3.4) extrapolates hyperparameter scaling laws to the 1.2 billion-parameter regime. Together, these protocols ensure principled, fair, and reproducible comparisons across different optimizers. We present the optimal configurations found and how loss changes with respect to each hyperparameter in Appendix˜B and hope that this can facilitate future research. We also open-source the code (https://github.com/marin-community/marin/tree/kaiyue/optimizers) and the corresponding WandB runs (https://wandb.ai/marin-community/optimizer-scaling).
3.1 General Experimental Setup
Following OLMo et al. (2025), we conduct all our experiments on a large-scale pretraining corpus composed of three publicly available datasets, tokenized with the Llama3 tokenizer: DCLM-baseline (3.8 trillion tokens, Li et al. (2025b)), StarCoder V2 Data (0.25 trillion tokens, Lozhkov et al. (2024)), and ProofPile 2 (55 billion tokens, Azerbayev et al. (2024)).
Optimizer
Acknowledgement
This work was supported by the Google TPU Research Cloud (TRC), the Stanford HAI–Google Cloud Credits Program, and NSF IIS 2211780, and is a part of the Marin Project. The authors thank Evan Walters, Omead Pooladzandi, Jiacheng You, and Zhiyuan Li for discussions and help during our research.
Appendix A Optimizer Definitions
In this section, we present the algorithm for each optimizer we evaluated.
Throughout this section, we use the following notation: for model parameters, for gradients at step , for learning rate, for weight decay, for moment decay rates, for numerical stability, for gradient norm, and for first and second moments. All operations are element-wise unless specified.
Based on Appendices˜C and D, we fitted our scaling law for the 1.2B run of Muon, NAdamW and AdamW, we round the fitted value to our grid of hyperparameter and used hyperparameters are shown in Appendix˜E.
After training the models, we fitted a scaling law for Muon and AdamW of the following form:
The fitted values are
For AdamW, , with a RMS error of .
For Muon , with a RMS error of .
This scaling law predicts that when the parameter scale reaches 7B, Muon will actually result in a higher loss compared to AdamW in 1 Chinchilla regime.
B.2 Sophia Experiments
We performed a Phase I experiment with Sophia, with detailed hyperparameter settings shown in Appendix˜C. We found that Sophia tends to underperform AdamW in smaller compute regimes and eventually slightly outperforms AdamW when either the model size or the data size increases.
Figure 7: Sophia Experiments. Left: Loss curve of Sophia and AdamW in Chinchilla setting. Right: Loss curve of Sophia and AdamW for 130M model size.
B.3 High Data-to-model Ratio
Figure 8: More Case Studies. Experiment with 130M Chinchilla setting, SOAP outperforms Muon in the overtraining setting.
B.4 Evaluation Performance
Table 6: Evaluation Performance for Mars, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.327
0.377
0.402
0.451
OPENBOOKQA
0.274
0.304
0.298
0.302
WINOGRANDE
0.521
0.512
0.526
0.522
PIQA
0.625
0.652
0.662
0.678
BOOLQ
0.616
0.504
0.569
0.557
WSC273
0.557
0.560
0.579
0.590
HELLASWAG 0SHOT
0.308
0.343
0.365
0.399
ARC CHALLENGE
0.225
0.258
0.272
0.247
ARC EASY
0.462
0.517
0.532
0.544
COPA
0.650
0.690
0.660
0.710
FINAL C4 LOSS
3.537
3.396
3.323
3.247
Table 7: Evaluation Performance for Mars, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.450
0.493
0.514
0.538
OPENBOOKQA
0.300
0.312
0.328
0.328
WINOGRANDE
0.534
0.557
0.549
0.561
PIQA
0.680
0.686
0.695
0.720
BOOLQ
0.480
0.533
0.612
0.608
WSC273
0.586
0.608
0.667
0.667
HELLASWAG 0SHOT
0.395
0.437
0.471
0.502
ARC CHALLENGE
0.253
0.269
0.298
0.307
ARC EASY
0.547
0.580
0.608
0.616
COPA
0.700
0.740
0.720
0.730
FINAL C4 LOSS
3.249
3.158
3.097
3.040
Table 8: Evaluation Performance for Mars, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.513
0.548
0.576
0.601
OPENBOOKQA
0.330
0.336
0.366
0.378
WINOGRANDE
0.548
0.559
0.586
0.601
PIQA
0.704
0.717
0.725
0.735
BOOLQ
0.559
0.598
0.625
0.631
WSC273
0.601
0.696
0.736
0.780
HELLASWAG 0SHOT
0.462
0.513
0.546
0.579
ARC CHALLENGE
0.294
0.317
0.333
0.351
ARC EASY
0.611
0.648
0.665
0.676
COPA
0.730
0.760
0.750
0.750
FINAL C4 LOSS
3.101
3.015
2.955
2.906
Table 9: Evaluation Performance for Muon, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.358
0.396
0.431
0.456
OPENBOOKQA
0.286
0.288
0.318
0.300
WINOGRANDE
0.515
0.518
0.530
0.541
PIQA
0.640
0.660
0.670
0.678
BOOLQ
0.571
0.504
0.583
0.452
WSC273
0.553
0.564
0.579
0.582
HELLASWAG 0SHOT
0.330
0.354
0.381
0.407
ARC CHALLENGE
0.231
0.244
0.269
0.276
ARC EASY
0.493
0.511
0.543
0.559
COPA
0.650
0.720
0.730
0.660
FINAL C4 LOSS
3.464
3.369
3.296
3.240
Table 10: Evaluation Performance for Muon, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.459
0.492
0.529
0.550
OPENBOOKQA
0.302
0.310
0.330
0.346
WINOGRANDE
0.504
0.534
0.548
0.570
PIQA
0.679
0.689
0.695
0.714
BOOLQ
0.482
0.555
0.599
0.570
WSC273
0.615
0.641
0.659
0.656
HELLASWAG 0SHOT
0.406
0.447
0.481
0.509
ARC CHALLENGE
0.270
0.289
0.290
0.316
ARC EASY
0.564
0.591
0.620
0.638
COPA
0.670
0.730
0.700
0.700
FINAL C4 LOSS
3.224
3.143
3.079
3.032
Table 11: Evaluation Performance for Muon, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.520
0.559
0.597
0.611
OPENBOOKQA
0.328
0.340
0.342
0.368
WINOGRANDE
0.560
0.567
0.594
0.590
PIQA
0.713
0.718
0.730
0.736
BOOLQ
0.590
0.555
0.633
0.613
WSC273
0.659
0.667
0.762
0.736
HELLASWAG 0SHOT
0.482
0.525
0.554
0.587
ARC CHALLENGE
0.300
0.318
0.351
0.348
ARC EASY
0.623
0.640
0.671
0.674
COPA
0.730
0.760
0.730
0.780
FINAL C4 LOSS
3.073
3.002
2.945
2.900
Table 12: Evaluation Performance for Lion, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.316
0.380
0.417
0.450
OPENBOOKQA
0.280
0.296
0.312
0.292
WINOGRANDE
0.503
0.518
0.519
0.521
PIQA
0.629
0.649
0.663
0.680
BOOLQ
0.603
0.507
0.419
0.489
WSC273
0.564
0.538
0.590
0.623
HELLASWAG 0SHOT
0.304
0.343
0.366
0.402
ARC CHALLENGE
0.229
0.272
0.262
0.272
ARC EASY
0.464
0.500
0.536
0.539
COPA
0.700
0.670
0.730
0.730
FINAL C4 LOSS
3.552
3.409
3.331
3.252
Table 13: Evaluation Performance for Lion, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.428
0.471
0.509
0.535
OPENBOOKQA
0.306
0.320
0.326
0.346
WINOGRANDE
0.506
0.530
0.532
0.564
PIQA
0.676
0.686
0.699
0.713
BOOLQ
0.486
0.490
0.609
0.593
WSC273
0.579
0.615
0.608
0.667
HELLASWAG 0SHOT
0.386
0.428
0.466
0.502
ARC CHALLENGE
0.267
0.296
0.294
0.322
ARC EASY
0.535
0.586
0.600
0.618
COPA
0.690
0.730
0.710
0.760
FINAL C4 LOSS
3.268
3.170
3.100
3.046
Table 14: Evaluation Performance for Lion, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.507
0.543
0.569
0.597
OPENBOOKQA
0.336
0.334
0.356
0.358
WINOGRANDE
0.560
0.548
0.581
0.590
PIQA
0.699
0.713
0.726
0.740
BOOLQ
0.600
0.546
0.612
0.631
WSC273
0.608
0.667
0.685
0.700
HELLASWAG 0SHOT
0.463
0.509
0.545
0.571
ARC CHALLENGE
0.303
0.320
0.340
0.371
ARC EASY
0.603
0.638
0.664
0.673
COPA
0.710
0.750
0.740
0.770
FINAL C4 LOSS
3.108
3.029
2.965
2.915
Table 15: Evaluation Performance for NAdamW, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.325
0.383
0.413
0.446
OPENBOOKQA
0.284
0.286
0.308
0.322
WINOGRANDE
0.515
0.507
0.546
0.507
PIQA
0.633
0.652
0.667
0.682
BOOLQ
0.591
0.569
0.459
0.557
WSC273
0.535
0.560
0.612
0.597
HELLASWAG 0SHOT
0.312
0.348
0.368
0.399
ARC CHALLENGE
0.239
0.245
0.242
0.270
ARC EASY
0.465
0.503
0.518
0.547
COPA
0.680
0.700
0.650
0.720
FINAL C4 LOSS
3.531
3.394
3.319
3.251
Table 16: Evaluation Performance for NAdamW, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.437
0.477
0.514
0.543
OPENBOOKQA
0.310
0.332
0.316
0.348
WINOGRANDE
0.520
0.536
0.549
0.564
PIQA
0.681
0.693
0.708
0.712
BOOLQ
0.539
0.564
0.520
0.612
WSC273
0.582
0.637
0.645
0.700
HELLASWAG 0SHOT
0.397
0.434
0.475
0.505
ARC CHALLENGE
0.264
0.275
0.292
0.307
ARC EASY
0.553
0.589
0.605
0.632
COPA
0.710
0.730
0.700
0.750
FINAL C4 LOSS
3.248
3.160
3.090
3.039
Table 17: Evaluation Performance for NAdamW, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.500
0.551
0.577
0.599
OPENBOOKQA
0.332
0.350
0.358
0.342
WINOGRANDE
0.553
0.571
0.594
0.597
PIQA
0.695
0.719
0.731
0.738
BOOLQ
0.613
0.606
0.639
0.618
WSC273
0.637
0.689
0.718
0.714
HELLASWAG 0SHOT
0.469
0.513
0.550
0.580
ARC CHALLENGE
0.294
0.322
0.342
0.354
ARC EASY
0.605
0.646
0.657
0.674
COPA
0.700
0.720
0.760
0.780
FINAL C4 LOSS
3.100
3.013
2.955
2.907
Table 18: Evaluation Performance for Kron, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.341
0.387
0.426
0.451
OPENBOOKQA
0.268
0.278
0.296
0.314
WINOGRANDE
0.520
0.522
0.515
0.543
PIQA
0.636
0.662
0.662
0.683
BOOLQ
0.533
0.530
0.575
0.557
WSC273
0.564
0.564
0.553
0.604
HELLASWAG 0SHOT
0.323
0.347
0.374
0.397
ARC CHALLENGE
0.235
0.259
0.256
0.281
ARC EASY
0.487
0.519
0.543
0.572
COPA
0.700
0.690
0.710
0.710
FINAL C4 LOSS
3.492
3.389
3.307
3.239
Table 19: Evaluation Performance for Kron, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.429
0.485
0.516
0.548
OPENBOOKQA
0.310
0.314
0.314
0.326
WINOGRANDE
0.527
0.534
0.542
0.571
PIQA
0.675
0.690
0.706
0.709
BOOLQ
0.487
0.475
0.610
0.611
WSC273
0.546
0.619
0.659
0.718
HELLASWAG 0SHOT
0.396
0.438
0.479
0.501
ARC CHALLENGE
0.262
0.285
0.298
0.323
ARC EASY
0.551
0.586
0.618
0.632
COPA
0.670
0.720
0.720
0.750
FINAL C4 LOSS
3.244
3.151
3.083
3.031
Table 20: Evaluation Performance for Kron, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.512
0.549
0.569
0.603
OPENBOOKQA
0.338
0.332
0.350
0.352
WINOGRANDE
0.536
0.567
0.575
0.601
PIQA
0.721
0.713
0.724
0.739
BOOLQ
0.529
0.560
0.537
0.559
WSC273
0.648
0.663
0.696
0.762
HELLASWAG 0SHOT
0.474
0.516
0.556
0.582
ARC CHALLENGE
0.287
0.312
0.344
0.353
ARC EASY
0.616
0.648
0.662
0.681
COPA
0.710
0.720
0.790
0.750
FINAL C4 LOSS
3.084
3.009
2.946
2.900
Table 21: Evaluation Performance for Scion, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.355
0.382
0.427
0.446
OPENBOOKQA
0.282
0.300
0.288
0.314
WINOGRANDE
0.515
0.502
0.502
0.535
PIQA
0.641
0.655
0.675
0.684
BOOLQ
0.574
0.518
0.532
0.474
WSC273
0.505
0.531
0.553
0.608
HELLASWAG 0SHOT
0.323
0.354
0.374
0.401
ARC CHALLENGE
0.247
0.241
0.276
0.264
ARC EASY
0.501
0.513
0.541
0.557
COPA
0.650
0.660
0.660
0.680
FINAL C4 LOSS
3.477
3.379
3.302
3.246
Table 22: Evaluation Performance for Scion, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.445
0.489
0.511
0.540
OPENBOOKQA
0.306
0.308
0.326
0.358
WINOGRANDE
0.531
0.550
0.555
0.566
PIQA
0.684
0.689
0.701
0.712
BOOLQ
0.576
0.610
0.574
0.585
WSC273
0.597
0.612
0.685
0.696
HELLASWAG 0SHOT
0.406
0.442
0.478
0.506
ARC CHALLENGE
0.272
0.285
0.298
0.316
ARC EASY
0.564
0.592
0.625
0.633
COPA
0.680
0.730
0.700
0.740
FINAL C4 LOSS
3.232
3.152
3.086
3.039
Table 23: Evaluation Performance for Scion, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.521
0.552
0.575
0.601
OPENBOOKQA
0.338
0.362
0.366
0.374
WINOGRANDE
0.569
0.551
0.605
0.586
PIQA
0.704
0.717
0.736
0.743
BOOLQ
0.575
0.615
0.620
0.641
WSC273
0.674
0.692
0.736
0.736
HELLASWAG 0SHOT
0.481
0.520
0.549
0.581
ARC CHALLENGE
0.289
0.314
0.346
0.354
ARC EASY
0.625
0.647
0.667
0.686
COPA
0.730
0.730
0.750
0.770
FINAL C4 LOSS
3.080
3.007
2.952
2.904
Table 24: Evaluation Performance for Cautious, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.314
0.385
0.420
0.458
OPENBOOKQA
0.298
0.280
0.308
0.318
WINOGRANDE
0.515
0.520
0.519
0.508
PIQA
0.638
0.652
0.664
0.669
BOOLQ
0.580
0.512
0.512
0.550
WSC273
0.560
0.527
0.593
0.604
HELLASWAG 0SHOT
0.314
0.347
0.371
0.401
ARC CHALLENGE
0.239
0.244
0.261
0.270
ARC EASY
0.476
0.508
0.532
0.544
COPA
0.690
0.720
0.710
0.690
FINAL C4 LOSS
3.535
3.403
3.334
3.253
Table 25: Evaluation Performance for Cautious, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.437
0.479
0.510
0.544
OPENBOOKQA
0.316
0.314
0.326
0.328
WINOGRANDE
0.515
0.533
0.530
0.576
PIQA
0.677
0.699
0.707
0.711
BOOLQ
0.547
0.492
0.584
0.574
WSC273
0.568
0.608
0.634
0.667
HELLASWAG 0SHOT
0.397
0.432
0.477
0.505
ARC CHALLENGE
0.267
0.280
0.292
0.308
ARC EASY
0.541
0.569
0.606
0.618
COPA
0.670
0.730
0.760
0.720
FINAL C4 LOSS
3.260
3.165
3.094
3.043
Table 26: Evaluation Performance for Cautious, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.509
0.549
0.578
0.604
OPENBOOKQA
0.318
0.338
0.342
0.364
WINOGRANDE
0.548
0.575
0.571
0.605
PIQA
0.705
0.715
0.733
0.739
BOOLQ
0.601
0.611
0.582
0.606
WSC273
0.619
0.718
0.659
0.736
HELLASWAG 0SHOT
0.467
0.514
0.550
0.583
ARC CHALLENGE
0.305
0.315
0.344
0.358
ARC EASY
0.609
0.649
0.676
0.689
COPA
0.730
0.760
0.770
0.770
FINAL C4 LOSS
3.100
3.017
2.956
2.910
Table 27: Evaluation Performance for SOAP, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.344
0.404
0.434
0.458
OPENBOOKQA
0.280
0.270
0.300
0.318
WINOGRANDE
0.511
0.510
0.523
0.517
PIQA
0.643
0.652
0.666
0.676
BOOLQ
0.569
0.491
0.519
0.559
WSC273
0.513
0.568
0.597
0.612
HELLASWAG 0SHOT
0.321
0.351
0.376
0.401
ARC CHALLENGE
0.247
0.255
0.270
0.270
ARC EASY
0.475
0.519
0.537
0.571
COPA
0.690
0.710
0.720
0.710
FINAL C4 LOSS
3.487
3.376
3.295
3.240
Table 28: Evaluation Performance for SOAP, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.001
0.492
0.518
0.541
OPENBOOKQA
0.256
0.324
0.324
0.334
WINOGRANDE
0.506
0.534
0.546
0.555
PIQA
0.546
0.689
0.706
0.711
BOOLQ
0.402
0.593
0.565
0.593
WSC273
0.498
0.619
0.703
0.670
HELLASWAG 0SHOT
0.257
0.443
0.478
0.506
ARC CHALLENGE
0.219
0.281
0.305
0.324
ARC EASY
0.309
0.590
0.612
0.632
COPA
0.590
0.740
0.720
0.700
FINAL C4 LOSS
5.437
3.147
3.082
3.030
Table 29: Evaluation Performance for SOAP, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.513
0.555
0.582
0.606
OPENBOOKQA
0.308
0.312
0.368
0.350
WINOGRANDE
0.548
0.562
0.594
0.583
PIQA
0.695
0.714
0.729
0.739
BOOLQ
0.563
0.623
0.596
0.643
WSC273
0.648
0.703
0.722
0.755
HELLASWAG 0SHOT
0.478
0.523
0.553
0.586
ARC CHALLENGE
0.292
0.345
0.334
0.356
ARC EASY
0.610
0.656
0.655
0.678
COPA
0.740
0.760
0.740
0.770
FINAL C4 LOSS
3.079
3.004
2.957
2.899
Table 30: Evaluation Performance for AdamW, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.322
0.380
0.410
0.442
OPENBOOKQA
0.306
0.290
0.298
0.292
WINOGRANDE
0.506
0.519
0.515
0.530
PIQA
0.632
0.663
0.665
0.674
BOOLQ
0.568
0.415
0.553
0.576
WSC273
0.582
0.553
0.564
0.623
HELLASWAG 0SHOT
0.312
0.340
0.365
0.391
ARC CHALLENGE
0.229
0.255
0.262
0.270
ARC EASY
0.463
0.497
0.523
0.544
COPA
0.660
0.650
0.720
0.730
FINAL C4 LOSS
3.529
3.409
3.322
3.262
Table 31: Evaluation Performance for AdamW, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.435
0.484
0.500
0.537
OPENBOOKQA
0.300
0.312
0.328
0.318
WINOGRANDE
0.519
0.530
0.542
0.552
PIQA
0.673
0.691
0.697
0.713
BOOLQ
0.454
0.496
0.508
0.602
WSC273
0.586
0.626
0.667
0.692
HELLASWAG 0SHOT
0.386
0.432
0.471
0.501
ARC CHALLENGE
0.269
0.295
0.304
0.310
ARC EASY
0.535
0.581
0.612
0.616
COPA
0.700
0.700
0.690
0.710
FINAL C4 LOSS
3.264
3.166
3.094
3.043
Table 32: Evaluation Performance for AdamW, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.500
0.540
0.570
0.591
OPENBOOKQA
0.320
0.336
0.342
0.356
WINOGRANDE
0.539
0.586
0.583
0.601
PIQA
0.705
0.719
0.732
0.739
BOOLQ
0.555
0.615
0.596
0.609
WSC273
0.604
0.656
0.703
0.729
HELLASWAG 0SHOT
0.456
0.507
0.543
0.578
ARC CHALLENGE
0.299
0.311
0.323
0.349
ARC EASY
0.613
0.639
0.669
0.688
COPA
0.740
0.680
0.710
0.780
FINAL C4 LOSS
3.110
3.023
2.958
2.913
Table 33: Evaluation Performance for Adam-Mini, Model Size = 130m
DATA SIZE
2B
5B
10B
21B
PERFORMANCE METRIC
LAMBADA OPENAI
0.317
0.369
0.412
0.444
OPENBOOKQA
0.286
0.282
0.282
0.302
WINOGRANDE
0.525
0.500
0.515
0.528
PIQA
0.640
0.655
0.662
0.670
BOOLQ
0.506
0.496
0.445
0.583
WSC273
0.571
0.557
0.546
0.634
HELLASWAG 0SHOT
0.310
0.337
0.364
0.390
ARC CHALLENGE
0.241
0.247
0.255
0.264
ARC EASY
0.465
0.495
0.525
0.555
COPA
0.690
0.710
0.730
0.700
FINAL C4 LOSS
3.542
3.416
3.328
3.266
Table 34: Evaluation Performance for Adam-Mini, Model Size = 300m
DATA SIZE
6B
12B
24B
48B
PERFORMANCE METRIC
LAMBADA OPENAI
0.429
0.468
0.509
0.535
OPENBOOKQA
0.300
0.318
0.318
0.350
WINOGRANDE
0.502
0.528
0.545
0.586
PIQA
0.688
0.693
0.683
0.701
BOOLQ
0.480
0.546
0.609
0.573
WSC273
0.586
0.546
0.593
0.689
HELLASWAG 0SHOT
0.381
0.424
0.464
0.492
ARC CHALLENGE
0.270
0.292
0.294
0.300
ARC EASY
0.545
0.579
0.607
0.625
COPA
0.700
0.720
0.740
0.730
FINAL C4 LOSS
3.272
3.178
3.103
3.049
Table 35: Evaluation Performance for Adam-Mini, Model Size = 520m
DATA SIZE
10B
21B
42B
85B
PERFORMANCE METRIC
LAMBADA OPENAI
0.500
0.538
0.576
0.604
OPENBOOKQA
0.306
0.326
0.330
0.368
WINOGRANDE
0.534
0.561
0.571
0.594
PIQA
0.707
0.721
0.723
0.733
BOOLQ
0.543
0.626
0.619
0.536
WSC273
0.619
0.700
0.733
0.744
HELLASWAG 0SHOT
0.459
0.505
0.541
0.576
ARC CHALLENGE
0.288
0.327
0.332
0.352
ARC EASY
0.613
0.638
0.654
0.671
COPA
0.710
0.720
0.760
0.760
FINAL C4 LOSS
3.112
3.027
2.966
2.912
Appendix C Hyperparameter Ablation in Phase I
We reported the results for the optimizers we swept in Phase I. The result is formulated as follows: the first row shows the approximately best configuration found and the following rows show the results for the 1-dimensional ablations centered around the found configuration. The loss presented here is the final loss on the C4/EN validation set.
C.1 Sweeping Results for AdamW
Table 36: Hyperparameter ablation for AdamW on 130m on 1x Chinchilla Data
This paper benchmarks the performance of 11 proposed optimizers and show vastly different speed-up ratio than prior works reported. In this section, we will compare the setup of our experiments with prior works with the hope of understanding the difference.
1.
Sophia Liu et al. [2024a] (2) This work utilizes a small peak learning rate of learning rate smaller than 6e-4 (similar to the one shown in Figure˜1 Top Left). The reason for the small peak learning rate is likely 2-fold: (i) the authors are training on a pretraining dataset PILE with lower quality compared to the current pretraining dataset and (ii) in the implementation of Levanter that the authors used, the data shuffling is not completely random and instead is correlated on every compute node. Upon reproducing the results, we note that this difference can significantly impact the stability of the training process and a complete random shuffling is crucial for the usage of a large learning rate.
2.
MARS Yuan et al. [2025] (2) This papers considers a similar setup as Sophia and uses a similar AdamW baseline. We note that in the first version of the paper, the authors also reported that increasing the learning rate of AdamW to 3e-3 can significantly improve the performance of AdamW (see Figure 6 of Yuan et al. [2025] arxiv version 1).
3.
Soap Vyas et al. [2025] (1.4) The actual speedup of Soap on 300M and 520M models are 1.2 to 1.3, which is only slightly lower than the claimed 1.4 speedup. We note that our implementation of Soap is slightly different from the one used in the paper that we performs blocking of weight in order to reduce the memory footprint and further uses bfloat16 for the momentum in the 1.2B experiments. Both modifications may lead to slightly lower step-wise performance.
4.
Muon Jordan et al. [2024], Liu et al. [2025a] (2) The speedup of Muon reported in different works are vastly different. In the original Nanogpt speedrun, Muon achieves 1.3 speedup over AdamW. Later the reproduction of Kimi reported a much higher speedup of 2. We note that the Kimi version utilizes a notably low learning rate for AdamW (8e-4 to 9e-4 for model between 400M to 1.5B) in the scaling experiments. We also note that the smaller learning rate is important for hyperparameter transfer from AdamW when roughly matching the update norm and AdamW can perform better higher learning rate in our experiments.
Further, their comparison of Muon and AdamW on the MoE experiments compare two models with not fully decayed learning rate and this may significantly favors Muon, as shown in Figure˜5.
It is later shown independently in the work of Essential AI AI et al. [2025] that Muon’s token efficiency compared to AdamW is only 1.1 to 1.2.
5.
Cautious Liang et al. [2025], Block-wise Learning Rate Adam Wang et al. [2025], FOCUS Liu et al. [2025c] report 2 speedup over AdamW. These papers use a similar baseline as Sophia and MARS.
6.
SWAN Ma et al. [2025] and DION Ahn et al. [2025] report 2-3 speedup over AdamW. The comparison between these two optimizers and AdamW is carried out on a smaller than 1 Chinchilla regime, where the speed-up of matrix-based optimizer may be larger. We also note that we didn’t consider the communication cost, which is the main focus of DION.
7.
SPlus Frans et al. [2025] reports a 2 speedup over AdamW. This work considers an atypical setup where the model is trained with a constant learning rate.